# Bringing Back the Context: Camera Trap Species Identification as Link Prediction on Multimodal Knowledge Graphs

### v3

@@ 5.4.3 Performance Comparison for Under-represented Species
The iWildCam2020-WILDS dataset exhibits a long-tail species distribution [30], posing challenges for accurately recognizing species that are under-represented in the training set. We compare the performance of our best-performing model (COSMO with location context) to the baseline Empirical Risk Minimization (ResNet-50) model (Table 6). We chose examples whose labels have a maximum of 100 instances in the training set and report the overall test set accuracy for this subset of species. This includes species like banded palm civet, Brazilian cottontail, and leopard which are listed as vulnerable in IUCN’s list of threatened species [63]. We observe that COSMO outperforms the ERM baseline by $2.7$%, which is a $16.6$% relative improvement. These findings illustrate the potential of our model to mitigate sample inefficiency in existing approaches for under-represented species by utilizing multimodal context information.
6 Discussion and Conclusion
@@
In this work, we presented a novel framework in which the species classification task is reformulated as link prediction in a multimodal KG of species images and their associated heterogeneous contexts. This enables a unified way to leverage various forms of multimodal context, e.g., numerical, categorical, and taxonomy information associated with images for species identification in camera traps. Through our experiments, we demonstrate that our framework achieves superior out-of-distribution (OOD) generalization and competitive performance with state-of-the-art for species classification on the iWildCam2020-WILDS and Snapshot Mountain Zebra datasets. Additionally, our framework demonstrates improved sample efficiency in recognizing under-represented and vulnerable wildlife species.
@@
We assume that there is a perfect linkage between these contexts and the corresponding images in the training set. However, in scenarios where such linkage is unavailable, the training procedure may introduce noise, which could lead to inferior generalization capabilities in the model. Additionally, it is important to note that the effectiveness of diverse contexts varies based on their informativeness for the given task. Interestingly, combining two or more contexts could degrade performance compared to using a single context type in some cases (Table 1). We hypothesize that specific metadata, like location, could exert a stronger regularization influence on improving generalization in species recognition tasks than other metadata. To improve this aspect, future work will involve enabling the model to assign higher importance to more informative metadata.
@@
Furthermore, we are interested in training a foundation model for camera trap species identification across a wider spectrum of species. This model is expected to generalize more effectively to novel camera trap deployments across the world. Additionally, we would like to focus on integrating a more comprehensive spectrum of diverse contexts such as temperature, weather conditions, habitat, and sequence information for use with real-world camera trap deployments.
@@ Acknowledgements
The authors would thank colleagues from the OSU NLP group for constructive feedback. This research was sponsored in part by NSF OAC 2112606, NSF OAC 2118240, Cisco, and the Ohio Supercomputer Center [14].
@@ Appendix A Datasets
The iWildCam2020-WILDS dataset [30] and Snapshot Mountain Zebra dataset [48] are released under the Community Data License Agreement (CDLA). The Open Tree Taxonomy [46] is licensed under Creative Commons Attribution 1.0 Generic license. Both licenses permit their use for academic research in their original form. The iWildCam2020-WILDS label space consists of 182 species. Out of these 182 species, 155 species have support in the OTT taxonomy.
Appendix B Hyperparameter and Experiment Details
The hidden dimension of the multimodal KG embedding model is set to 512, with a batch size of 16. In our experiments, iWildCam2020-WILDS was trained for 12 epochs, whereas Snapshot Mountain Zebra required 15 epochs for optimal performance. All images are resized to $448\times 448$. We use Adam [29] optimizer with a learning rate of 3e-5 and 1e-3 for the image encoder and the rest of the parameters, respectively. For the location and time attributes, we use a 3-layer MLP that projects the feature input dimension to the embedding dimension and uses PReLU as the activation function. We use early stopping based on validation accuracy to prevent overfitting. The early stopping patience parameter is set to 5 epochs. Table B.1 and Table B.2 show the training time of COSMO under different settings for iWildCam2020-WILDS and Snapshot Mountain Zebra, respectively. We performed all experiments using a single Nvidia RTX A6000 GPU.
@@ Appendix B Hyperparameter and Experiment Details
Table B.1: Training time for different ablations of COSMO (iWildCam2020-WILDS). All experiments use a single Nvidia RTX A6000 GPU.
@@
Table B.2: Training time for different ablations of COSMO (Snapshot Mountain Zebra). All experiments use a single Nvidia RTX A6000 GPU.
@@ Appendix C Time Attribute Analysis
Figure C.1: Species probabilities conditioned on day/night for the 10 most frequent species in the training set (iWildCam2020-WILDS). Animal species demonstrate distinct temporal preferences for their daily activities, as evidenced by the contrasting probabilities observed during day and night.

### v5

@@ 5.4.3. Performance Comparison for Under-represented Species
The iWildCam2020-WILDS dataset exhibits a long-tail species distribution (Koh et al., 2021), posing challenges for accurately recognizing species that are under-represented in the training set. We compare the performance of our best-performing model (COSMO with location context) to the baseline Empirical Risk Minimization (ResNet-50) model (Table 6). We focus on examples whose labels have a maximum of 100 instances in the training set and report the overall test set accuracy for this subset of species. This selection includes species like banded palm civet, Brazilian cottontail, and leopard, all classified as vulnerable in IUCN’s list of threatened species (Vié et al., 2009). We observe that COSMO outperforms the ERM baseline by $2.7$%, which is a $16.6$% relative improvement. These findings illustrate the potential of our model to mitigate sample inefficiency in existing approaches for under-represented species by utilizing multimodal context information.
6. Discussion and Conclusion
@@
In this work, we presented a novel framework in which the species classification task is reformulated as link prediction in a multimodal KG of species images and their diverse contextual information. This enables a unified way to leverage various forms of multimodal context, e.g., numerical, categorical, and taxonomy information associated with images for species classification in camera traps. Through our experiments, we demonstrate that our framework achieves superior out-of-distribution generalization and competitive performance with state-of-the-art for species classification on the iWildCam2020-WILDS and Snapshot Mountain Zebra datasets. Additionally, our framework exhibits improved sample efficiency in recognizing under-represented and vulnerable wildlife species.
@@
We assume that there is a perfect linkage between these contexts and the corresponding images in the training set. However, in scenarios where such linkage is unavailable, the training procedure may introduce noise, which could lead to inferior generalization capabilities in the model. Additionally, it is important to note that the effectiveness of diverse contexts varies based on their informativeness for the given task. Interestingly, combining two or more contexts could degrade performance compared to using a single context type in some cases (Table 1). We posit that specific metadata, like location, might have a stronger regularization effect on improving generalization in species recognition tasks than other metadata. To address this, future work will involve enabling the model to assign greater importance to more informative metadata.
@@
Furthermore, we are interested in training a foundation model for camera trap species classification across a wider spectrum of species. This model should demonstrate enhanced generalization capabilities for new camera trap setups worldwide. Additionally, we aim to integrate a broader spectrum of diverse contexts such as temperature, weather conditions, habitat, and sequence information for use with real-world camera trap deployments.
@@ Acknowledgements
The authors would like to thank colleagues from the OSU NLP group for their constructive feedback. This research was sponsored in part by NSF OAC 2112606, NSF OAC 2118240, Cisco, and the Ohio Supercomputer Center (Center, 1987).
@@ Appendix A Datasets
The iWildCam2020-WILDS dataset (Koh et al., 2021) and Snapshot Mountain Zebra dataset (Pardo et al., 2021) are released under the Community Data License Agreement (CDLA). The Open Tree Taxonomy (OpenTreeofLife et al., 2019) is licensed under Creative Commons Attribution 1.0 Generic license. Both licenses permit their use for academic research in their original form. The iWildCam2020-WILDS label space consists of 182 species. Out of these 182 species, 155 species have support in the OTT taxonomy.
Appendix B Experiment Details
Table B.1 and Table B.2 show the training time of COSMO under different settings for iWildCam2020-WILDS and Snapshot Mountain Zebra, respectively. We performed all experiments using a single Nvidia RTX A6000 GPU.
Table B.1. Training time for different ablations of COSMO (iWildCam2020-WILDS). All experiments use a single Nvidia RTX A6000 GPU.
@@ Appendix B Experiment Details
Table B.2. Training time for different ablations of COSMO (Snapshot Mountain Zebra). All experiments use a single Nvidia RTX A6000 GPU.
@@ Appendix C Time Attribute Analysis
Figure C.1. Species probabilities conditioned on day/night for the 10 most frequent species in the training set (iWildCam2020-WILDS). Animal species demonstrate distinct temporal preferences for their daily activities, as evidenced by the contrasting probabilities observed during day and night.